OpenStudy (anonymous):

What is the solution to the equation log4 x + log4 2 = 3 ?

OpenStudy (anonymous):

log4 (2x) =3 2x=4^3 2x=64 x=32

OpenStudy (jackellyn):

\[\log _{a}x=N , \ a ^{N}=x\]

OpenStudy (anonymous):

log A +logB =log(AB)